OpenStudy (anonymous):

***PLEASE! Help Needed!!?!*** ***Medal AND Fan Rewarded*** Find b, given that a = 18.2, Angle B = 62°, and Angle C = 48°. Round answers to the nearest whole number.

OpenStudy (raden):

m<A + m<B + <C = 180 can u find the angle A, first ?

OpenStudy (anonymous):

58

OpenStudy (raden):

m<A +62° + 48° = 180 so, <A = ... ?

OpenStudy (anonymous):

58...

OpenStudy (raden):

still incorrect m<A +62° + 48° = 180° m<A + 110° = 180° m<A = 180° - 110° m<A = 70°

OpenStudy (raden):

now, to get b use the sine rule : b sinA = a sinB b sin 70° = 18.2 sin 62° solve for b